OpenStudy (anonymous):

find the equation of the tangent line to the curve y = ln(x2 − 9x + 1), (9, 0)

myininaya (myininaya):

Need help finding y'? That is the first step. Hint. (ln(u))'=u'/u

OpenStudy (anonymous):

what do you after that step

myininaya (myininaya):

equation of the tangent line can be found using the point-slope form for a line y-y1=m(x-x1) m=f'(9)

myininaya (myininaya):

Your point is (9,0)